OpenStudy (anonymous):

can you show me the steps to this problem -8(4+9x) = 7(-2 - 11x)

OpenStudy (anonymous):

-32-72x=-14-77x -32+14=-77x+72x

OpenStudy (anonymous):

solve for x

jhonyy9 (jhonyy9):

-32-72x= -14-77x 5x=18 x=18/5